OpenStudy (jmark):

Average of 50 numbers is 38. If 45 and 55 are removed sum gets decreased by 100. Find their average.

Parth (parthkohli):

\[38 = \frac{S}{50}\Rightarrow S = 50\cdot 38 = 1900\]Now\[S' = 1900 - 55 - 45=1800\]\[N' = 50 - 2 = 48\]\[A' = \frac{S'}{N'}\]